i am stuck on the following problem

suppose h is a function defined by h(x)=(fog)(x). if h(x)=absolute value -11x-6, what is f(x)

any help would be appreciated

Answer by richard1234(7193)

There are many answers. Assuming you mean h(x) = |-11x-6|, you could have g(x) = -11x-6, f(x) = |x|, for example. Or g(x) = |-11x-6| and f(x) = x.